# COMPAS: A Distributed Multi-Party SWAP Test for Parallel Quantum Algorithms

### v1

@@ 1. Introduction
In distributed quantum computing, efficient execution of algorithms relies on high-rate and low-latency communication between modules (parekh_quantum_2021; Cacciapuoti2020quantum). The basic resource for such a modular system is a Bell pair shared between physical qubits in different nodes, enabling quantum-state teleportation and remote gate operations. Developing an interface with sufficient rate and fidelity to support high-fidelity computation remains a significant challenge. There has been tremendous experimental progress developing interconnections among quantum systems, such as using cryogenic microwave links superconducting qubits or optical photons for atomic qubits (ang2024arquin; main2025distributed). Yet, state-of-the-art results for remote entanglement between neutral atom or trapped ion qubits only achieves rates of a few hundred Hz with entangling fidelities in the high 90% (Stephenson_2020; jing2019entanglement; Ritter_2012; Hucul_2014; Pompili_2021). Entanglement distillation protocols have also been developed to further boost the fidelity of entanglement while maintaining a high rate and low latency (rozpkedek2018optimizing). Despite this progress, architectures that can efficiently connect modular and distributed quantum hardware with generic high-level quantum algorithms are much underdeveloped.

### v2

@@ Abstract.
Quantum computing; Distributed architectures; Entanglement distillation
isbn: 978-1-4503-XXXX-X/2018/06journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790143isbn: 979-8-4007-2359-9/2026/03ccs: Hardware Quantum computationccs: Computer systems organization Distributed architectures
@@ 1. Introduction
In distributed quantum computing, efficient execution of algorithms relies on high-rate and low-latency communication between modules (parekh_quantum_2021; Cacciapuoti2020quantum). The basic resource for such a modular system is a Bell pair shared between physical qubits in different nodes, enabling quantum-state teleportation and remote gate operations. Developing an interface with sufficient rate and fidelity to support high-fidelity computation remains a significant challenge. There has been tremendous experimental progress developing interconnections among quantum systems, such as using cryogenic microwave links for superconducting qubits or optical photons for atomic qubits (ang2024arquin; main2025distributed). Yet, state-of-the-art results for remote entanglement between neutral atom or trapped ion qubits only achieves rates of a few hundred Hz with entangling fidelities in the high 90% (Stephenson_2020; jing2019entanglement; Ritter_2012; Hucul_2014; Pompili_2021). Entanglement distillation protocols have also been developed to further boost the fidelity of entanglement while maintaining a high rate and low latency (rozpkedek2018optimizing). Despite this progress, architectures that can efficiently connect modular and distributed quantum hardware with generic high-level quantum algorithms are much underdeveloped.
@@ 2. Background
2.1. High-Level Overview
[…]
The multi-party SWAP test is a standard tool for comparing many quantum states. It operates by extracting the trace of a product of states, a quantity in the interval $[0,1]$ that provides information about the states’ overlaps, spectra, and entanglement structure. However, implementing the multi-party SWAP test on near-term hardware is expensive due to hardware limitations. In spite of this, distributed quantum computing offers a potential remedy by distributing computational tasks across a network of interconnected quantum processing units (QPUs). This reduces the computational load on each device, but can incur additional communication overhead.
In this work, we present a distributed implementation of the multi-party SWAP test that simultaneously achieves optimal latency and memory usage. To our knowledge, this is the first approach to do so. Our construction, and distributed quantum protocols in general, use pairs of correlated (entangled) states known as Bell pairs, which are distributed between parties in advance. These pre-shared resources enable spatially separated QPUs to interact with each other without directly transmitting quantum states during runtime. We leverage this capability to distribute the SWAP test over many parties while consuming Bell pairs as a resource.